two boat each move a velocity v relative to the water and both cross a river of breadth a running with uniform velocity V.they start together one boat crossing by shortest path and the otherin shortest time.find the difference between the time of arrival.
1
Expert's answer
2020-04-23T18:58:28-0400

Shortest path

Relative Velocity of boat is in a direction at an angle "\theta" with the vertical

"vsin\theta=V"

"cos\theta=\sqrt{\frac{v^2-V^2}{v^2}}"

"Time=\frac{Distance}{speed}"

"t_1=d/(vcos\theta)"

"t_1=\frac{d}{\sqrt{v^2-V^2}}"

Shortest time

Relative velocity of boat is perpendicular to the velocity of river.

"t_2=d\/v"

"t_2-t_1=d[\frac1 v-\frac{1}{\sqrt{v^2-V^2}}]"